OpenStudy (anonymous):

What is the sum? -9 + 3

OpenStudy (alexandervonhumboldt2):

-9+3=-6 is this a real question?

TheSmartOne (thesmartone):

-9 + 3 = 3 - 9 = ?

OpenStudy (thread252):

-6

OpenStudy (alexandervonhumboldt2):

too mahy people, thus it is spam idiotic logic isn't it?

rebeccaxhawaii (rebeccaxhawaii):

-6

OpenStudy (chantysquirrel1129**):

Maybe she's in like 5th grade :p I helped a 5th grader on here once lol ....

OpenStudy (bibby):

god damn 5th grades are getting THICC

OpenStudy (chantysquirrel1129**):

I'm serious lol. I was like what are you doing on here??? Who knows xp

OpenStudy (thread252):

the answer is B